solve using the pythagorem theorem c=? a=10,b=24
Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12844)

c=? a=10,b=24
a^2+b^2=c^2
10^2+24^2=c^2
100+576=c^2
c^2=675
c=sqrt675
c=26 ans.

Answer by solver91311(24713)



Presuming that , , and have their traditional meanings, that is that and are the legs of the right triangle and is the hypotenuse:



So for your problem



Now just do the arithmetic. Note that 5-12-13 is a Pythagorean Triple and that 10 is 2 times 5 and 24 is 2 times 12. Therefore if you get an answer for that is different from 26, go back and check your arithmetic.

John